# Spectral selectors and contact orderability

### v1

@@ 4.1. The spectral metrics
Corollary 4.1 ([58, Theorems 2.17 and 2.26]).
On either $\mathcal{L}$ or $-~\widetilde{\mathcal{L}}$ orderable, one has
@@
for any supporting contact form $\alpha$. Moreover, $\mathrm{d}_{\mathrm{spec}}^{\alpha}(\Lambda,\Lambda^{\prime})=0$ implies $\Lambda=\Lambda^{\prime}$ on $\mathcal{L}$ (resp. $\Pi\Lambda=\Pi\Lambda^{\prime}$ on $-~\widetilde{\mathcal{L}}$) while $\gamma^{\alpha}(\Lambda,\Lambda^{\prime})=0$ implies $\Lambda=\phi^{\alpha}_{t}\Lambda^{\prime}$ for some $t\in\mathbb{R}$ on $\mathcal{L}$ (resp. $\Pi\Lambda=\phi^{\alpha}_{t}\Pi\Lambda^{\prime}$ on $-~\widetilde{\mathcal{L}}$).
@@
The triangular inequality satisfied by $\ell_{+}^{\alpha}$ implies that the maps ±$\ell_{\pm}^{\alpha}(\cdot,\Lambda_{0})$ are $\mathrm{d}_{\mathrm{spec}}^{\alpha}$-1-Lipschitz:
@@
Let us recall that a partially ordered metric space $(Z,d,\leq)$ is a metric space $(Z,d)$ endowed with a partial order $\leq$ such that $a\leq b\leq c$ implies $d(a,b)\leq d(a,c)$. The following statement directly follows from the definitions.
@@
Let $O$ be either $\mathcal{L}$, $-~\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$, associated to some contact manifold $(M,\ker\alpha)$ (and possibly some closed Legendrian submanifold of $M$). If $O$ is orderable, then $(O,\mathrm{d}_{\mathrm{spec}}^{\alpha},\preceq)$ is a partially ordered metric space.
@@
Proposition 4.4 ([58, Proposition 2.3]).
Let $O$ be either $\mathcal{L}$, $-~\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$, associated to some contact manifold $(M,\ker\alpha)$ (and possibly some closed Legendrian submanifold of $M$) and let us assume that $O$ is orderable. The topology induced by the spectral distance $\mathrm{d}_{\mathrm{spec}}^{\alpha}$ is the interval topology, that is the topology generated by the basis of open subsets
@@
Let $(M,\xi)$ be a contact manifold and $\alpha$ be a complete contact form supporting $\xi$. Let $f:M\to\mathbb{R}$ be a bounded map and $\beta:=e^{f}\alpha$. Then, on either $\mathcal{L}$, $-~\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$ orderable for which selectors are well-defined,
@@
One could ask whether one could have more: a natural invariant metric (bi-invariant in the case of $\mathcal{G}$ and $\widetilde{\mathcal{G}}$). According to Fraser-Polterovich-Rosen [35, Theorem 3.1], such metric would essentially be discrete because of the possibility to squeeze a Darboux ball of $(M,\xi)$ into an arbitrarily small open subset. Therefore, the induced topology will not be that interesting, although discrete bi-invariant metrics could carry interesting information, especially regarding their asymptotic behaviors (see Sections 4.3, 4.4 and 4.5 below).
@@ 4.2. Spectrally robust Legendrian interlinkings
In particular the hypothesis of Theorem 4.7 is satisfied whenever $\Lambda_{0}$ and $\Lambda_{1}$ are two different elements in $\widetilde{\mathcal{L}}$ satisfying $\Lambda_{0}\preceq\Lambda_{1}$ thanks to Lemma 3.6. Moreover for any $\varphi\in\widetilde{\mathcal{G}}$ the sign invariance guarantees that $\ell_{+}^{\alpha}(\Lambda_{1},\Lambda_{0})>0$ if and only if $\ell_{+}^{\alpha}\left(\varphi(\Lambda_{1}),\varphi(\Lambda_{0})\right)>0$ (Lemma 3.2). Therefore, Theorem 4.7 generalizes [32, Theorem 1.5 (i)].
@@
The spectral-robustness now follows from the $1$-$\mathrm{d}_{\mathrm{spec}}^{\alpha}$-Lipchitzness of +$\ell^{\alpha}_{+}(\Lambda_{1},\cdot)$ and +$\ell^{\alpha}_{+}(\cdot,\Lambda_{0})$ (by (14) and the Poincaré duality property). It implies Hofer and $C^{1}$-robustness by Corollary 4.1. ∎
@@ 4.3. Unboundedness of Hofer type pseudo-metrics
Let $\Lambda_{*}\subset(M,\xi)$ be a closed Legendrian submanifold of a contact manifold such that $\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$) is orderable and $\alpha$ be a complete contact form supporting $\xi$. Then $\mathrm{d}_{\mathrm{SCH}}^{\alpha}(\Lambda,\Lambda^{\prime})=0$ if and only if $\Lambda=\Lambda^{\prime}$ (resp. $\Pi\Lambda=\Pi\Lambda^{\prime}$), for any $\Lambda,\Lambda^{\prime}\in\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$). Moreover, for all $\Lambda\in\mathcal{L}$ (resp. $-~\widetilde{\mathcal{L}}$), $(\phi_{t}^{\alpha}\Lambda)_{t\in\mathbb{R}}$ is a geodesic for $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$:
@@
The non-degeneracy of $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$ when $\mathcal{L}$ is orderable was originally due to Hedicke [41, Theorem 5.2], while the non-degeneracy of $\mathrm{d}_{\mathrm{SH}}^{\alpha}$ was proven by Shelukhin without any condition on $\mathcal{G}$ [68], as mentioned earlier in Section 2.3. The unboundedness of $\mathrm{d}_{\mathrm{SH}}^{\alpha}$ and $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$ had already been proven by Hedicke in the special case where $(M,\xi)$ is a unit tangent bundle with open cover and $\Lambda_{*}$ is a fiber of the bundle (cf Example 2.10.5) [41, Theorems 5.7 and 5.8]. For the open contact manifold $(\mathbb{R}^{2n}\times S^{1},\ker\alpha_{\mathrm{st}})$ defined in Example 2.10.3 (here $N:=\mathbb{R}^{n}$ and $\alpha_{\mathrm{st}}:=\mathrm{d}z-\lambda$), some geodesics of $\mathrm{d}_{\mathrm{SH}}^{\alpha_{\mathrm{st}}}$ have also been characterized in [7] using, among other things, the spectrality of Sandon contact selectors [64].
@@
Let us set $c:=c_{+}^{\alpha}(\varphi)$. This implies that $\varphi\preceq\phi_{c+\varepsilon}^{\alpha}$ for any $\varepsilon>0$ and so $r_{\pm}(p,\varphi)\leq r_{\pm}(p,\phi_{c+\varepsilon}^{\alpha})$. Since $\phi_{c+\varepsilon}^{\alpha}$ is generated by the constant function equal to $c+\varepsilon$ by the result of Eliashberg and Polterovich discussed above $r_{\pm}(p,\phi_{c+\varepsilon}^{\alpha})=c+\varepsilon$. Letting $\varepsilon$ going to 00 we get the inequality $r_{\pm}(p,\varphi)\leq c_{+}^{\alpha}(\varphi)$. The proof for the other inequality follows the same lines. ∎
@@
is a well defined conjugation invariant norm, i.e. whether it takes finite values or not. This question was already asked by Shelukhin [68, Question 18] for the Hofer-Shelukhin norm $\left|\cdot\right|_{\mathrm{SH}}^{\alpha}$. We do not know the answer to this question, however if we extend the action by conjugation to the whole group of coorientation preserving contactomorphisms that are not necessarily isotopic to the identity we have the following proposition.
@@
    It is interesting to note that if $(M,\ker\alpha)$ is a closed contact manifold such that the Reeb flow of $\alpha$ is $1$-periodic and $\widetilde{\mathcal{G}}(M)$ is orderable then
@@ 4.4. Colin-Sandon discriminant and oscillation metrics
Let $\Lambda_{*}$ be a closed Legendrian submanifold of a cooriented contact manifold $(M,\xi)$ (not necessarily closed) and let $\mathcal{L}:=\mathcal{L}(\Lambda_{*})$ (resp. ~:=~$\widetilde{\mathcal{L}}:=\widetilde{\mathcal{L}}(\Lambda_{*})$). In this specific section, given $\Lambda_{0},\Lambda_{1}\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$), by a path $\gamma:\Lambda_{0}\leadsto\Lambda_{1}$ we will mean a continuous map $\gamma:[0,1]\to\mathcal{L}$ (resp. $[0,1]\to\widetilde{\mathcal{L}}$) such that $\gamma(i)=\Lambda_{i}$ for $i\in\{0,1\}$. The concatenation $\gamma_{1}\cdot\gamma_{2}$ of two paths $\gamma_{1}:\Lambda_{0}\leadsto\Lambda_{1}$ and $\gamma_{2}:\Lambda_{1}\leadsto\Lambda_{2}$ in $\mathcal{L}$ (resp. in $\widetilde{\mathcal{L}}$) is by definition the path $\Lambda_{0}\leadsto\Lambda_{2}$,
@@
The reverse path $\bar{\gamma}$ of a path $\gamma:\Lambda_{0}\leadsto\Lambda_{1}$ is the path $\Lambda_{1}\leadsto\Lambda_{0}$, $t\mapsto\gamma(1-t)$. By definition of the topology defined on $\mathcal{L}$, for any path $\gamma:[0,1]\to\mathcal{L}$, there exists a continuous map $j_{\gamma}:[0,1]\times\Lambda_{*}\to M$ such that for every $t\in[0,1]$, $j_{\gamma}(t,\cdot)$ is a diffeomorphism between $\Lambda_{*}$ and $\gamma(t)$. A path $\gamma$ in $\mathcal{L}$ will be called embedded if there exists a continuous map $j_{\gamma}$ as above that is a smooth embedding $[0,1]\times\Lambda_{*}\hookrightarrow M$. A path $\gamma$ in $\widetilde{\mathcal{L}}$ will be called embedded if $\Pi\gamma:t\mapsto\Pi(\gamma(t))$ is an embedded path.
In [26, Section 8], Colin and Sandon defined the discriminant length of a path $\gamma$ in $\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$) as the integral number
@@
with convention $\ell_{\mathrm{disc}}(\gamma)=0$ if $\gamma$ is a constant map. Colin-Sandon proved that $\ell_{\mathrm{disc}}$ takes values in $\mathbb{N}$ (with convention $0\in\mathbb{N}$) and that it induces an integral metric on $\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$) called the discriminant metric and defined by
@@
In [26], the unboundedness of these two distances was only proven in the case of ~$\widetilde{\mathcal{L}}(\mathbb{R}\mathrm{P}^{n})$ (see Example 2.10.1), some $1$-dimensional Legendrian knots, and ~$\widetilde{\mathcal{L}}(p(0_{N}))$ defined in Example 2.10.3.
@@
Let us assume there exists a contact form $\alpha$ supporting $\xi$ the Reeb flow of which is $T$-periodic for some $T>0$. If $-~\widetilde{\mathcal{L}}$ is orderable, then
@@
By definition, there exists embedded paths $\gamma_{i}:\Lambda_{i-1}\leadsto\Lambda_{i}$, $1\leq i\leq n$, with $\Lambda_{n}=\Lambda$. Let us consider the maps $f_{i}:t\mapsto\ell_{+}^{\alpha}(\gamma_{i}(t),\gamma_{i}(0))$, $1\leq i\leq n$. By the normalization property of $\ell_{+}^{\alpha}$, $f_{i}(0)=0$ for all $i$. By continuity of $\ell_{+}^{\alpha}$, the $f_{i}$’s are continuous maps. By spectrality of $\ell_{+}^{\alpha}$, if $f_{i}(t)=T$ for some $t\in(0,1]$ and $1\leq i\leq n$, it would imply that $\Pi(\gamma_{i}(t))$ intersects $\phi^{\alpha}_{T}\Pi(\gamma_{i}(0))$. But $\phi^{\alpha}_{T}=\mathrm{id}$ by assumption so $\gamma_{i}$ would not be an embedded path. Therefore, $f_{i}$ does not take the value $T$. By continuity, it implies that the $f_{i}$’s take their values in $(-T,T)$. Now, by the triangular inequality,
@@
For the second inequality, one can apply the same strategy. Let $k:=\mathrm{d}_{\mathrm{CS,osc}}^{+}(\Lambda_{0},\Lambda)$. By definition, there exists embedded monotone paths $\gamma_{i}:\Lambda_{i-1}\leadsto\Lambda_{i}$, $1\leq i\leq n$ for some $n\in\mathbb{N}$ such that $k$ of them are positive, $n-k$ of them are negative and $\Lambda_{n}=\Lambda$. By the triangular inequality,
@@
Let us assume that there exists a contact form $\alpha$ supporting $\xi$ the Reeb flow of which is $T$-periodic for some $T>0$. If Λ~normal-$-~\Lambda\in\widetilde{\mathcal{L}}$ orderable is such that $\phi_{t}^{\alpha}\Pi\Lambda\cap\Pi\Lambda=\emptyset$ for all $t\notin T\mathbb{Z}$, then the isotopy $(\phi^{\alpha}_{tT}\Lambda)_{t\in\mathbb{R}}$ defines a geodesic of both the discriminant and the Colin-Sandon oscillation metrics, in the sense that

### v2

@@ 4.1. The spectral metrics
Corollary 4.1 ([59, Theorems 2.17 and 2.26]).
On either $\mathcal{L}$ or $\widetilde{\mathcal{L}}$ orderable, one has
@@
for any supporting contact form $\alpha$. Moreover, $\mathrm{d}_{\mathrm{spec}}^{\alpha}(\Lambda,\Lambda^{\prime})=0$ implies $\Lambda=\Lambda^{\prime}$ on $\mathcal{L}$ (resp. $\Pi\Lambda=\Pi\Lambda^{\prime}$ on $\widetilde{\mathcal{L}}$) while $\gamma^{\alpha}(\Lambda,\Lambda^{\prime})=0$ implies $\Lambda=\phi^{\alpha}_{t}\Lambda^{\prime}$ for some $t\in\mathbb{R}$ on $\mathcal{L}$ (resp. $\Pi\Lambda=\phi^{\alpha}_{t}\Pi\Lambda^{\prime}$ on $\widetilde{\mathcal{L}}$).
@@
The triangle inequality satisfied by $\ell_{+}^{\alpha}$ implies that the maps ±$\ell_{\pm}^{\alpha}(\cdot,\Lambda_{0})$ are $\mathrm{d}_{\mathrm{spec}}^{\alpha}$-1-Lipschitz:
@@
[…]
@@
Let $O$ be either $\mathcal{L}$, $\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$, associated to some contact manifold $(M,\ker\alpha)$ (and possibly some closed Legendrian submanifold of $M$). If $O$ is orderable, then $(O,\mathrm{d}_{\mathrm{spec}}^{\alpha},\preceq)$ is a partially ordered metric space.
@@
Proposition 4.4 ([59, Proposition 2.3]).
Let $O$ be either $\mathcal{L}$, $\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$, associated to some contact manifold $(M,\ker\alpha)$ (and possibly some closed Legendrian submanifold of $M$) and let us assume that $O$ is orderable. The topology induced by the spectral distance $\mathrm{d}_{\mathrm{spec}}^{\alpha}$ is the interval topology, that is the topology generated by the basis of open subsets
@@
Let $(M,\xi)$ be a contact manifold and $\alpha$ be a complete contact form supporting $\xi$. Let $f:M\to\mathbb{R}$ be a bounded map and $\beta:=e^{f}\alpha$. Then, on either $\mathcal{L}$, $\widetilde{\mathcal{L}}$, $\mathcal{G}$ or $\widetilde{\mathcal{G}}$ orderable for which selectors are well-defined,
@@
One could ask whether one could have more: a natural invariant metric (bi-invariant in the case of $\mathcal{G}$ and $\widetilde{\mathcal{G}}$). According to Fraser-Polterovich-Rosen [36, Theorem 3.1], such metric would essentially be discrete because of the possibility to squeeze a Darboux ball of $(M,\xi)$ into an arbitrarily small open subset. Therefore, the induced topology will not be that interesting, although discrete bi-invariant metrics could carry interesting information, especially regarding their asymptotic behaviors (see Sections 4.3, 4.4 and 4.5 below).
@@ 4.2. Spectrally robust Legendrian interlinkings
In particular the hypothesis of Theorem 4.7 is satisfied whenever $\Lambda_{0}$ and $\Lambda_{1}$ are two different elements in $\widetilde{\mathcal{L}}$ satisfying $\Lambda_{0}\preceq\Lambda_{1}$ thanks to Lemma 3.6. Moreover for any $\varphi\in\widetilde{\mathcal{G}}$ the sign invariance guarantees that $\ell_{+}^{\alpha}(\Lambda_{1},\Lambda_{0})>0$ if and only if $\ell_{+}^{\alpha}\left(\varphi(\Lambda_{1}),\varphi(\Lambda_{0})\right)>0$ (Lemma 3.2). Therefore, Theorem 4.7 generalizes [33, Theorem 1.5 (i)] since $\widetilde{\mathcal{L}}$ of the zero-section of the 111-jet bundle $J^{1}N$ is orderable for any closed manifold $N$ (see the third point in Example 2.14).
@@
[…]
@@ 4.3. Unboundedness of Hofer type pseudo-metrics
Let $\Lambda_{*}\subset(M,\xi)$ be a closed Legendrian submanifold of a contact manifold such that $\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$) is orderable and $\alpha$ be a complete contact form supporting $\xi$. Then $\mathrm{d}_{\mathrm{SCH}}^{\alpha}(\Lambda,\Lambda^{\prime})=0$ if and only if $\Lambda=\Lambda^{\prime}$ (resp. $\Pi\Lambda=\Pi\Lambda^{\prime}$), for any $\Lambda,\Lambda^{\prime}\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$). Moreover, for all $\Lambda\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$), $(\phi_{t}^{\alpha}\Lambda)_{t\in\mathbb{R}}$ is a geodesic for $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$:
@@
The non-degeneracy of $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$ when $\mathcal{L}$ is orderable was originally due to Hedicke [42, Theorem 5.2], while the non-degeneracy of $\mathrm{d}_{\mathrm{SH}}^{\alpha}$ was proven by Shelukhin without any condition on $\mathcal{G}$ [70], as mentioned earlier in Section 2.3. The unboundedness of $\mathrm{d}_{\mathrm{SH}}^{\alpha}$ and $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$ had already been proven by Hedicke in the special case where $(M,\xi)$ is a unit tangent bundle with open cover and $\Lambda_{*}$ is a fiber of the bundle (cf Example 2.14.5) [42, Theorems 5.7 and 5.8]. For the open contact manifold $(\mathbb{R}^{2n}\times S^{1},\ker\alpha_{\mathrm{st}})$ defined in Example 2.14.3 (here $N:=\mathbb{R}^{n}$ and $\alpha_{\mathrm{st}}:=\mathrm{d}z-\lambda$), some geodesics of $\mathrm{d}_{\mathrm{SH}}^{\alpha_{\mathrm{st}}}$ have also been characterized in [8] using, among other things, the spectrality of Sandon contact selectors [65].
@@
Let us set $c:=c_{+}^{\alpha}(\varphi)$. This implies that $\varphi\preceq\phi_{c+\varepsilon}^{\alpha}$ for any $\varepsilon>0$ and so $r_{\pm}(p,\varphi)\leq r_{\pm}(p,\phi_{c+\varepsilon}^{\alpha})$. Since $\phi_{c+\varepsilon}^{\alpha}$ is generated by the constant function equal to $c+\varepsilon$ by the result of Eliashberg and Polterovich discussed above $r_{\pm}(p,\phi_{c+\varepsilon}^{\alpha})=c+\varepsilon$. Letting $\varepsilon$ going to 0 we get the inequality $r_{\pm}(p,\varphi)\leq c_{+}^{\alpha}(\varphi)$. The proof for the other inequality follows the same lines. ∎
@@
is a well defined conjugation invariant norm, i.e. whether it takes finite values or not. This question was already asked by Shelukhin [70, Question 18] for the Hofer-Shelukhin norm $\left|\cdot\right|_{\mathrm{SH}}^{\alpha}$. We do not know the answer to this question, however if we extend the action by conjugation to the whole group of coorientation preserving contactomorphisms that are not necessarily isotopic to the identity we have the following proposition.
@@
    It is interesting to note that if $(M,\ker\alpha)$ is a closed contact manifold such that the Reeb flow of $\alpha$ is 111-periodic and $\widetilde{\mathcal{G}}(M)$ is orderable then
@@ 4.4. Colin-Sandon discriminant and oscillation metrics
Let $\Lambda_{*}$ be a closed Legendrian submanifold of a cooriented contact manifold $(M,\xi)$ (not necessarily closed) and let $\mathcal{L}:=\mathcal{L}(\Lambda_{*})$ (resp. ~:=~$\widetilde{\mathcal{L}}:=\widetilde{\mathcal{L}}(\Lambda_{*})$). In this specific section, given $\Lambda_{0},\Lambda_{1}\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$), by a path $\gamma:\Lambda_{0}\leadsto\Lambda_{1}$ we will mean a $C^{1}$-continuous map $\gamma:[0,1]\to\mathcal{L}$ (resp. $[0,1]\to\widetilde{\mathcal{L}}$) such that $\gamma(i)=\Lambda_{i}$ for $i\in\{0,1\}$. The concatenation $\gamma_{1}\cdot\gamma_{2}$ of two paths $\gamma_{1}:\Lambda_{0}\leadsto\Lambda_{1}$ and $\gamma_{2}:\Lambda_{1}\leadsto\Lambda_{2}$ in $\mathcal{L}$ (resp. in $\widetilde{\mathcal{L}}$) is by definition the path $\Lambda_{0}\leadsto\Lambda_{2}$,
@@
The reverse path $\bar{\gamma}$ of a path $\gamma:\Lambda_{0}\leadsto\Lambda_{1}$ is the path $\Lambda_{1}\leadsto\Lambda_{0}$, $t\mapsto\gamma(1-t)$. By definition of the topology defined on $\mathcal{L}$, for any path $\gamma:[0,1]\to\mathcal{L}$, there exists a continuous map $j_{\gamma}:[0,1]\times\Lambda_{*}\to M$ such that for every $t\in[0,1]$, $j_{\gamma}(t,\cdot)$ is a diffeomorphism between $\Lambda_{*}$ and $\gamma(t)$. A path $\gamma$ in $\mathcal{L}$ will be called embedded if there exists a smooth map $j_{\gamma}$ as above that is a smooth embedding $[0,1]\times\Lambda_{*}\hookrightarrow M$. A path $\gamma$ in $\widetilde{\mathcal{L}}$ will be called embedded if $\Pi\gamma:t\mapsto\Pi(\gamma(t))$ is an embedded path.
In [27, Section 8], Colin and Sandon defined the discriminant length of a path $\gamma$ in $\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$) as the integral number
@@
with convention $\ell_{\mathrm{disc}}(\gamma)=0$ if $\gamma$ is a constant map and $\min\emptyset=+\infty$. Colin-Sandon proved that $\ell_{\mathrm{disc}}$ takes values in $\mathbb{N}$ (with convention $0\in\mathbb{N}$) and that it induces an integral metric on $\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$) called the discriminant metric and defined by
@@
In [27], the unboundedness of these two distances was only proven in the case of ~$\widetilde{\mathcal{L}}(\mathbb{R}\mathrm{P}^{n})$ (see Example 2.14.1), some 111-dimensional Legendrian knots, and ~$\widetilde{\mathcal{L}}(p(0_{N}))$ defined in Example 2.14.3.
@@
Let us assume there exists a contact form $\alpha$ supporting $\xi$ the Reeb flow of which is $T$-periodic for some $T>0$. If $\widetilde{\mathcal{L}}$ is orderable, then
@@
[…]
@@
For the second inequality, one can apply the same strategy. Let $k:=\mathrm{d}_{\mathrm{CS,osc}}^{+}(\Lambda_{0},\Lambda)$. By definition, there exists embedded monotone paths $\gamma_{i}:\Lambda_{i-1}\leadsto\Lambda_{i}$, $1\leq i\leq n$ for some $n\in\mathbb{N}$ such that $k$ of them are positive, $n-k$ of them are negative and $\Lambda_{n}=\Lambda$. By the triangle inequality,
@@
Let us assume that there exists a contact form $\alpha$ supporting $\xi$ the Reeb flow of which is $T$-periodic for some $T>0$. If $\Lambda\in\widetilde{\mathcal{L}}$ orderable is such that $\phi_{t}^{\alpha}\Pi\Lambda\cap\Pi\Lambda=\emptyset$ for all $t\notin T\mathbb{Z}$, then the isotopy $(\phi^{\alpha}_{tT}\Lambda)_{t\in\mathbb{R}}$ defines a geodesic of both the discriminant and the Colin-Sandon oscillation metrics, in the sense that